OpenStudy (otonogold):

What is the solution of the following system? {-3x-2y=-12 {9x+6y=-9

OpenStudy (anonymous):

Do u mean simultaneous equations?

OpenStudy (otonogold):

(2,1) (-2,-1) no solutions many solutions

OpenStudy (anonymous):

What is your mean ??? Is the first one a system ???

OpenStudy (otonogold):

yes, the first one is a system.

OpenStudy (anonymous):

And the secound is the answers ?

OpenStudy (otonogold):

The second problem? We are talking about these {-3x-2y=-12 {9x+6y=-9 right?

OpenStudy (anonymous):

http://www.webmath.com/solver2.html @OtonoGold

OpenStudy (anonymous):

K ! Look : -3x-2y=-12 9x+6y=-9 We have *3 for first : 3*(-3x-2y=-12) -9x-6y=-36 K ?!:)

OpenStudy (otonogold):

It dosent work @UditKulka

OpenStudy (otonogold):

so it will be many solutions? @E.ali

OpenStudy (anonymous):

Add both the equations

OpenStudy (anonymous):

Yes we have many solutions but my solutions was nt end !:)

OpenStudy (anonymous):

@E.ali is correct Add the equations and -9x & +9X Cancels out

OpenStudy (anonymous):

Thank you !:) You re best ! @UditKulka

OpenStudy (anonymous):

So there are no sollutions for this problem

OpenStudy (anonymous):

Welcome @E.ali

OpenStudy (anonymous):

Did u get it @OtonoGold